OpenStudy (anonymous):

Please help medals will be given!!! A grocery store sells pears for $1.99 per pound and apples for $1.49 per pound. A customer purchases 9 pounds of pears and apples for $15.41. How many pounds of apples did the customer buy?

OpenStudy (studygurl14):

Okay, so you know that a customer buys apples and pears and that the cost (without tax) adds up to 15.41. You don't know how many pounds of apples he bought. However, you do know how much each costs, and how many pounds of pears he bought. You can use this information and make an equation. A = pounds of apples P = pounds of pears A(1.49) + P(1.99) = 15.41 A(149) + P(199) = 1541 A(149) + 9(199) = 1541 Can you do the rest?

OpenStudy (studygurl14):

@maria.alison

OpenStudy (anonymous):

I think so....

OpenStudy (anonymous):

I got 0.0057745850805107

OpenStudy (anonymous):

I don't think that's right....

OpenStudy (studygurl14):

neither do I. Hm... for some reason this isn't coming out right. Are you sure you typed the question up right?

OpenStudy (anonymous):

Yeah, I copy pasted it

OpenStudy (anonymous):

The 9 pounds is apples and pears combined, not just pears. equation for cost like said A(149) + P(199) = 1541 equation for pounds purchased A+P=9 now you've got a system of equations. Know how to start?

OpenStudy (studygurl14):

hm.... Because I got a negative, which is impossible

OpenStudy (studygurl14):

Oh, ops. I read the question wrong....

OpenStudy (studygurl14):

Sorry!

OpenStudy (studygurl14):

What @CalebBeavers said.

OpenStudy (anonymous):

so what do I do then

OpenStudy (studygurl14):

Okay, so you have these two equations. A(149) + P(199) = 1541 A + P = 9 You will want to use substitution. And since the question is asking for how many pounds of apples, I'd get rid of the pears part. A + P = 9 --> P = 9 - A Now, plug in (9 - A) into the first equation. A(149) + (9 - A)(199) = 1541 Now what do you do?

OpenStudy (anonymous):

IDK

OpenStudy (studygurl14):

You would multiply (9 - A) and 199. Can you do that?

OpenStudy (anonymous):

171-9A

OpenStudy (studygurl14):

Um, no. 199( 9 - A) 199(9) - 199A What is 199 times 9 ?

OpenStudy (anonymous):

1791 sorry i typed it wrong the first time.

OpenStudy (studygurl14):

Good, now put that into the equation. 149A +199(9-A) = 1541 149A + 1791 - 199A = 1541 What next?

OpenStudy (anonymous):

subtract 1791 from both sides?

OpenStudy (studygurl14):

Sure, yeah. You could do that. If you did that, what would the equation be?

OpenStudy (anonymous):

149A-199A=-250

OpenStudy (anonymous):

So the answer is 5....

OpenStudy (anonymous):

gave you a medal

OpenStudy (anonymous):

peace out

OpenStudy (studygurl14):

good job!

OpenStudy (studygurl14):

and thanks

OpenStudy (studygurl14):

sorry I didn't continue doing the problem with you. computer troubles.

OpenStudy (anonymous):

Its fine, that last part was pretty simple

OpenStudy (studygurl14):

yep. anyway, the answer is 5

OpenStudy (anonymous):

k